I have a math project coming up and I have to find the solution to 2 equations. I'm really confused because both of the equations begin with 7= and I can't find x. My equations are y=11.25x-80 and y=7.50x-50. Can you please solve it and give me the steps so I understand. Thanks
Answer by mananth(16949)

y=11.25x-80 and y=7.50x-50
equate the y
11.25x-80=7.50x-50
11.25x-7.50x=80-50
3.75x=30
x=30/3.75
x= 8

plug x in any of the above equations to get y
y=10